OpenStudy (anonymous):

Ben wants to double his money in eight years by investing it in a savings account with interest compounded annually. What interest rate must the savings account offer in order for his money to double in eight years?

OpenStudy (anonymous):

The equation for compound interest is Where A is the amount accumulated P is the principle invested r is the interest rate n is the how often the interest is compounded each year t is the number of years A = P(1+r/n)^(n*t) A = 2P 2P = P(1+r/n)^(n*t) 2 = (1+r/n)^(n*t) 2 = (1+r)^(t) [n = 1, because the money is compounded annually] 2 = (1+r)^(8) [t = 8, because Ben wants the money to double in 8 years] 1+r = 8th root of 2 r = (8th root of 2) - 1 The rate should be about 9.05%. Hope you understand this problems :D

OpenStudy (anonymous):

Thank you so much. (:

OpenStudy (anonymous):

welcome :)